1,3,5,7,9,11,13,15,17,19 from these numbers, give any 5 numbers that total 50.
Answer by Edwin McCravy(20064)

There is no solution, because:

the sum of an odd number of odd numbers is always odd.

All those are odd numbers and 5 is an odd number of them,
but 50 is an even number, so there is no use to try.

You can get 49 with 5 of those numbers this way 3+5+7+15+19 = 49

You can get 51 with 5 of those numbers this way 5+7+11+13+15 = 51

With 6 of those numbers you can get 50 this way 3+5+7+9+11+15 = 50

But there is no way possible to get 5 odd numbers that total 50!!!

Edwin